OpenStudy (anonymous):

Giuliana has 22 quarters and dollar coins worth a total of $10.75. Which equation can be used to find d, the number of dollar coins Giuliana has? d – 22 + 0.25d = 10.75 0.25d + 22 – d = 10.75 0.25(22 – d) + d = 10.75 d + 0.25(d – 22) = 10.75

OpenStudy (anonymous):

Well, quarters (q) are worth $0.25 and dollar coins (d) are worth $1.00. You want to find out how many d and how many q that Giuliana has, and you know all her money adds up to 10.75. Let q equal the number of quarters and d equal the number of dollar coins. There's a total of 22 of both so, q + d = 22 and as for the price .25q + 1d = 10.75 Solve for q in the first equation by subtracting d from both sides and you get q = 22 - d Now substitute (22 - d) into the second equation for q and you get .25(22-d) + d = 10.75 So, it would appear that your answer is C.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

ok did it help you

OpenStudy (anonymous):

yes